OpenStudy (hotpinkgirl):

The ramp shown below is used to move crates of apples to loading docks at different heights. When the horizontal distance AB is 12 feet, the height of the loading dock, BC, is 6 feet. What is the height of the loading dock DE? 10 feet 12 feet 17 feet 18 feet

OpenStudy (hotpinkgirl):

@Cresciez

OpenStudy (cresciez):

Sorry, I'm not sure about this one :(

OpenStudy (hotpinkgirl):

thats ok

OpenStudy (michele_laino):

is BD=22?

OpenStudy (hotpinkgirl):

yes

OpenStudy (michele_laino):

since the triangles \(ABC\) and \(ADE\) are similar, then we can write this proportion: \[\huge 12:6 = \left( {22 + 12} \right):DE\]

OpenStudy (michele_laino):

If we apply the fundamental property of proportions, we can write this: \[\huge DE = \frac{{6 \cdot \left( {22 + 12} \right)}}{{12}} = ...?\]

OpenStudy (hotpinkgirl):

oh, is it 17 ?

OpenStudy (michele_laino):

that's right!

OpenStudy (hotpinkgirl):

k, thanks : )

OpenStudy (michele_laino):

:)